OpenStudy (anonymous):

if f(x)=2x^2+x-5, g(x)=2x-5, h(x)=x-2/1+x, find f^1(x)

OpenStudy (aum):

Are you asked to find the inverse function of f(x) ?

OpenStudy (anonymous):

yes

OpenStudy (aum):

f(x)=2x^2+x-5 y = 2x^2 + x - 5 y = 2(x^2 + 1/2x - 5/2) complete the square: y = 2{ (x+1/4)^2 - 1/16 - 5/2 } y / 2 = (x+1/4)^2 - 1/16 - 40/16 = (x+1/4)^2 - 41/16 y/2 + 41/16 = (x+1/4)^2 (8y + 41) / 16 = (x+1/4)^2 Take square root on both sides: sqrt(8y+41) / 4 = x + 1/4 x = { sqrt(8y+41) - 1 } / 4 interchange x and y: y = { sqrt(8x+41) - 1 } / 4

OpenStudy (anonymous):

Thank you!!!:)

OpenStudy (aum):

\[ f^{-1}(x) = \frac{\pm\sqrt{8x+41} - 1}{4} \]

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

What about h^-1(x)?

OpenStudy (anonymous):

Is it the same steps?

OpenStudy (aum):

To find the inverse of h(x): 1. Replace h(x) by y. 2. Solve for x in terms of y. 3. Interchange x and y 4. Replace y by h^(-1)(x)